# Fair Resource Allocation in Weakly Coupled Markov Decision Processes

### v1

@@ 3.2 The Count Aggregation MDP
Assuming symmetry across all $N$ sub-MDPs and using a permutation-invariant policy within a utilitarian framework allows us to simplify the global MDP by aggregating the sub-MDPs based on their state counts and tracking the number of actions taken in each state. Since each sub-MDP follows the same transition probabilities and reward structure, we can represent the entire system more compactly.
@@
Both representations lead to the same optimization problem as established in Gast et al., (2024) when the objective is utilitarian. Using the count representation, the mean expected total discounted reward $\bar{V}_{0}^{{\pi}_{{\bm{1}}/N}}$ for a WCMDP $\mathcal{M}^{(N)}$ with permutation invariant distribution $\mu^{(N)}$ and equal weights ${\bm{w}}_{{\bm{1}}/N}$ (Theorem 3.4) is then equivalent to the expected total discounted mean reward $\bar{V}_{0}^{{\pi}_{\phi}}$ for the count aggregation MDP $\mathcal{M}_{\phi}$ given the policy ${\pi}_{\phi}:\mathcal{S}^{(N)}_{f}\rightarrow\Delta(\mathcal{A}^{(N)}_{g_{s}})$ under aggregation mapping with initial distribution $\mu^{(N)}_{f}$, i.e., $\bar{V}_{0}^{{\pi}_{{\bm{1}}/N}}=\frac{1}{N}\sum_{n=1}^{N}V^{{\pi}_{{\bm{1}}/N}}_{0,n}=\frac{1}{S}\sum_{s=1}^{S}V^{{\pi}_{\phi}}_{0,s}=\bar{V}_{0}^{{\pi}_{\phi}}$. The objective in Eq. 2 is therefore reformulated as $\underset{{\pi}_{\phi}}{\max}\,\bar{V}_{0}^{{\pi}_{\phi}}$, i.e.,
@@
A LP method is provided to solve the count aggregation MDP in Appendix D.2.
@@ 4 Count-Proportion-Based DRL
We now consider the situation where the transition dynamics $p^{(N)}_{\phi}$ are unknown and the learner computes the (sub-)optimal policy through trial-and-error interactions with the environment. In Section 4.1, we introduce a count-proportion-based deep RL (CP-DRL) approach. This method incorporates a stochastic policy neural network with fixed-sized inputs and outputs, designed for optimizing resource allocation among stakeholders under constraints with count representation. In Section 4.2, we detail the priority based sampling procedure used to generate count actions.
@@ 4.1 Stochastic Policy Neural Network
One key property of the count aggregation MDP is that the dimensions of state space $\mathcal{S}^{(N)}_{f}$ and action space $\mathcal{A}^{(N)}_{g_{s}}$ are constant and irrespective of the number of sub-MDPs. To further simplify the analysis and eliminate the influence of $N$, we define the count state proportion as $\bar{\bm{x}}=\bm{x}/N$ and the resource proportion constraint for each resource $k$ as $\bar{b}_{k}=b_{k}/(N\max_{a\in\mathcal{A}}d_{k}(a))\in[0,1]$.
@@
Figure 1: CP-based stochastic policy neural network.
@@ 4.2 Priority-based Sampling Procedure
  Apply masking to ${\bm{U}}$ and update the forbidden set by ${\mathcal{F}}\leftarrow\mathcal{F}\cup\{(s,a)\mid u^{p}_{s,a}=0\text{ for }s\in[S]\text{ and }a\in[A]\}.$
@@
        Add $(s,a)$ to forbidden set ${\mathcal{F}}\leftarrow\mathcal{F}\cup\{(s,a)\}.$
@@ 5 Experimental Results
We apply our methods to the machine replacement problem (Delage and Mannor,, 2010; Akbarzadeh and Mahajan,, 2019), providing a scalable framework for evaluating the CP-DRL approach as problem size and complexity increase. We focus on a single resource ($K=1$) and binary action ($A=2$) for each machine, allowing validation against the Whittle index policy for RMABs (Whittle,, 1988). We applied various DRL algorithms, including Soft Actor Critic (SAC), Advantage Actor Critic (A2C), and Proximal Policy Optimization (PPO). Among these, PPO algorithm (Schulman et al.,, 2017) consistently delivers the most stable and high-quality performance. We thus choose PPO as the main algorithm for our CP-DRL approach (see Section 4). Our code is provided on Github222 https://anonymous.4open.science/r/GGF-WCMDP.
Machine Replacement Problem   The problem consists of $N$ identical machines following Markovian deterioration rules with $S$ states representing aging stages. The state space $\mathcal{S}^{(N)}$ is a Cartesian product. At each decision stage, actions $\bm{a}$ are applied to all machines under the resource constraint, with action $a_{n}$ representing operation (passive action 0) or replace (active action 1). Resource consumption $d_{n}(a_{n})$ is 1 for replacements and 0 for operation, with up to $b$ replacements per time step. Costs range from 0 to 1, transformed to fit the reward representation by multiplying by -1 and adding 1. Machines degrade if not replaced and remain in state $S$ until replaced. Refer to Appendix F.1 for cost structures and the transition probabilities. We choose the operational and replacement costs across two presets to capture different scenarios (see Appendix F.2 for details): i) Exponential-RCCC and ii) Quadratic-RCCC.
The goal is to find a fair policy that maximizes the GGF score over expected total discounted mean rewards with count aggregation MDP. In cases like electricity or telecommunication networks, where equipment is regionally distributed, a fair policy guarantees an equitable operation and replacement, thereby preventing frequent failures in specific areas that lead to unsatisfactory and unfair results for certain customers.
Experimental Setup   We designed a series of experiments to test the GGF-optimality, flexibility, scalability, and efficiency of our CP-DRL algorithm. We compare against optimal solutions (OPT) from the GGF-LP model (3) for small instances solved with Gurobi 10.0.3, the Whittle index policy (WIP) for RMABs, and a random (RDM) agent that selects actions randomly at each time step and averages the results over 10 independent runs.
@@
Experiment 2 (Flexibility)   The fixed-size input-output design allows CP-DRL to leverage multi-task training (MT) with varying machine numbers and resources. We refer to this multi-task extension as CP-DRL(MT). We trained the CP-DRL(MT) with $N\in\{2,3,4,5\}$, randomly switching configurations at the end of each episode over 2000 training episodes. CP-DRL(MT) was evaluated separately, and GGF values for WIP and RDM policies were obtained from 1000 Monte Carlo runs. The numbers following the plus-minus sign ($\pm$) represent the variance across 5 experiments with different random seeds in Table 2 and 2. Variances for WIP and RDM are minimal and omitted, with bold font indicating the best GGF scores at each row excluding optimal values. As shown in Table 2, CP-DRL(MT) consistently achieves scores very close to the OPT values as the number of machines increases from 2 to 4. For the 5-machine case, CP-DRL(MT) shows slightly better performance than the single-task CP-DRL. In Table 2, the single- and multi-task CP-DRL agents show slight variations in performance across different machine numbers. For $N=5$, CP-DRL achieves the best GGF score, slightly outperforming WIP.
Table 1: GGF Scores (Exponential-RCCC).
$N$ OPT WIP CP-DRL CP-DRL(MT) RDM
2 14.19 14.07 14.12 $\pm$ 0.01 14.11 $\pm$ 0.01 9.67
3 14.08 13.75 13.95 $\pm$ 0.02 13.89 $\pm$ 0.14 10.13
4 13.94 13.27 13.64 $\pm$ 0.05 13.59 $\pm$ 0.10 9.74
5 13.77 12.47 12.96 $\pm$ 0.01 13.28 $\pm$ 0.03 8.95
Table 2: GGF Scores (Quadratic-RCCC).
$N$ OPT WIP CP-DRL CP-DRL(MT) RDM
2 16.17 16.17 16.14 $\pm$ 0.00 16.14 $\pm$ 0.00 10.15
3 16.10 16.09 16.05 $\pm$ 0.00 16.05 $\pm$ 0.00 11.83
4 16.01 16.01 15.94 $\pm$ 0.00 15.94 $\pm$ 0.00 12.17
5 15.91 15.86 15.87 $\pm$ 0.02 15.86 $\pm$ 0.02 11.98
Experiment 3 (Scalability)   We assess CP-DRL scalability by increasing the number of machines while keeping the resource proportion at 0.1 for Exponential-RCCC instances. We refer to this scaled extension as CP-DRL(SC). Machine numbers vary from 10 to 100 to evaluate CP-DRL performance as the problem size grows. We also use CP-DRL(SC), trained on 10 machines with 1 unit of resource, and scale it to tasks with 20 to 100 machines. Figure 3(a) shows CP-DRL and CP-DRL(SC) consistently achieve higher GGF values than WIP as machine numbers increase. CP-DRL(SC) delivers results comparable to separately trained CP-DRL, reducing training time while maintaining similar performance. Both WIP and CP-DRL show linear growth in time consumption per episode as machine numbers scale up.
(a) GGF values for the number of
machines $N\in[10,100]$.
(b) Time per episode in seconds
with a resource ratio $b/N=0.1$.
(c) Time per episode in seconds
with a resource ratio $b/N=0.5$.
@@
Figure 3: (Colored) Scalability and Time Efficiency of CP-DRL. Subfigures (a) and (b) show the scalability of CP-DRL with a fixed resource ratio of 0.1. Subfigure (a) presents GGF values across different machine counts, with intervals representing standard deviation over 5 runs. Subfigure (b) and (c) depicts time per episode in seconds for a fixed resource ratio of 0.1 and 0.5, respectively. In all time plots, the green line represents WIP during MC evaluation, the blue line shows CP-DRL during training, and the orange line represents CP-DRL during MC evaluation.
@@
Experiment 4 (Efficiency)   In the GGF-LP model (3), the number of constraints grows exponentially with the number of machines $N$ as $N^{2}+S^{N}$, and the variables increase by $2N+(N+1)\cdot S^{N}$. Using the Count dual LP model (18) reduces the model size, but constraints still grow as $\binom{N+S-1}{S-1}$ and variables increase by $\binom{N+S-1}{S-1}\cdot A$. These growth patterns create computational challenges as problem size increases. Detailed time analysis for GGF-LP and Count dual LP models with $N$ from 2 to 7 is provided in Appendix F.4. In addition to the time per episode for a fixed ratio of 0.1 in Figure 3a, we analyze performance with a 0.5 ratio (Figure 3c) and varying machine proportions, keeping the number of machines fixed at 10. We evaluate CP-DRL over machine proportions from 0.1 to 0.9. The results show that time per episode increases linearly with machine count, while the training and evaluation times remain relatively stable. This indicates that the sampling procedure for legal actions is the primary bottleneck. Meanwhile, the resource ratio has minimal impact on computing times.
@@ 6 Conclusion
We incorporate the fairness consideration in terms of the generalized Gini function within the weakly coupled Markov decision processes, and define the GGF-WCMDP optimization problem. First, we present an exact method based on linear programming for solving it. We then derive an equivalent problem based on a utilitarian reduction when the WCMDP is symmetric, and show that the set of optimal permutation invariant policy for the utilitarian objective is also optimal for the original GGF-WCMDP problem. We further leverage this result by utilizing a count state representation and introduce a count-proportion-based deep RL approach to devise more efficient and scalable solutions. Our empirical results show that the proposed method using PPO as the RL algorithm consistently achieves high-quality GGF solutions. Moreover, the flexibility provided by the count-proportion approach offers possibilities for scaling up to more complex tasks and context where Whittle index policies are unavailable due to the violation of the indexibility property by the sub-MDPs.
@@ Appendix A Related Work
Fairness-aware learning is increasingly integrated into the decision-making ecosystem to accommodate minority interests. However, naively imposing fairness constraints can actually exacerbate inequity (Wen et al.,, 2021) if the feedback effects of the decisions are ignored. Many real-world fairness applications are not one-time static decisions (Zhao and Gordon,, 2019) and can thus be better modeled with sequential decision problems, which still remain relatively under-studied.
@@
Fairness with dynamics There are a few studies investigating fairness-aware sequential decision making. For instance, Liu et al., (2018) consider one-step delayed feedback effects, Creager et al., (2020) propose causal modeling of dynamical systems to address fairness, and Zhang et al., (2019) construct a user participation dynamics model where individuals respond to perceived decisions by leaving the system uniformly at random. These studies extend the fairness definition in temporally extended decision-making settings, but do not take feedback and learning into account that the system may fail to adapt to changing conditions. Alamdari et al., (2023) address this gap by introducing multi-stakeholder fairness as non-Markovian sequential decision-making and developing a Q-learning based algorithm with counterfactual experiences to enhance sample-efficient fair policy learning.
@@
Fairness in Markov decision processesZhang et al., (2020) consider how algorithmic decisions impact the evolution of feature space of the underlying population modeled as Markov decision processes (MDP) but is limited to binary decisions. Ghalme et al., 2022b study a fair resource allocation problem in the average MDP setting and proposes an approximate algorithm to compute the policy with sample complexity bounds. However, their definition of fairness is restricted to the minimum visitation frequency across all states, potentially resulting in an unbalanced rewards among users. Wen et al., (2021) develop fair decision-making policies in discounted MDPs, but the performance guarantees are achieved only under a loose condition. In contrast, our work takes into account a more comprehensive definition of fairness. Segal et al., (2023) investigate the impact of societal bias dynamics on long-term fairness and the interplay between utility and fairness under various optimization parameters. Additionally, Hassanzadeh et al., (2023) address a fair resource allocation problem similar to our work but in continuous state and action space. They define fairness to the agents considering all their allocations over the horizon under the Nash Social Welfare objective in hindsight.
@@
Fairness in reinforcement learningJabbari et al., (2017) initiate the meritocratic fairness notion from the multi-arm bandits setting to the reinforcement learning (RL) setting. Later, fairness consideration has been integrated in reinforcement learning to achieve fair solutions in different domains, including a fair vaccine allocation policy that equalizes outcomes in the population (Atwood et al.,, 2019), balancing between fairness and accuracy for interactive user recommendation (Liu et al.,, 2020; Ge et al.,, 2022), and fair IoT that continuously monitors the human state and changes in the environment to adapt its behavior accordingly (Elmalaki,, 2021). However, most work focuses on the impartiality aspect of fairness. Jiang and Lu, (2019) investigate multi-agent RL where fairness is defined over agents and encoded with a different welfare function, but the focus is on learning decentralized policies in a distributed way. We refer readers to two literature review papers by Gajane et al., (2022) and Reuel and Ma, (2024) on fairness considerations in RL, which provide comprehensive insights into current trends, challenges, and methodologies in the field.

### v2

@@ 3.2 The Count Aggregation MDP
Assuming symmetry across all $N$ sub-MDPs and using a permutation-invariant policy within a utilitarian framework allows us to simplify the global MDP by aggregating the sub-MDPs based on their state counts and tracking the number of actions taken in each state. Since each sub-MDP follows the same transition probabilities and reward structure, we can represent the entire system more compactly. This symmetry consideration is practical in many real-world applications where a large number of identical or interchangeable identities demand fair and efficient treatment, such as patients in healthcare or taxi drivers in public transportation services. By leveraging symmetry, we can reduce computational complexity for scalable fair solutions while inherently enforcing fairness as the policy treats all sub-MDPs equivalently.
@@
Both representations lead to the same optimization problem as established in Gast et al., (2024) when the objective is utilitarian. Using the count representation, the mean expected total discounted reward $\bar{V}_{0}^{{\pi}_{{\bm{1}}/N}}$ for a WCMDP $\mathcal{M}^{(N)}$ with permutation invariant distribution $\mu^{(N)}$ and equal weights ${\bm{w}}_{{\bm{1}}/N}$ (Theorem 3.4) is then equivalent to the expected total discounted mean reward $\bar{V}_{0}^{{\pi}_{\phi}}$ for the count aggregation MDP $\mathcal{M}_{\phi}$ given the policy ${\pi}_{\phi}:\mathcal{S}^{(N)}_{f}\rightarrow\Delta(\mathcal{A}^{(N)}_{g_{s}})$ under aggregation mapping with initial distribution $\mu^{(N)}_{f}$, i.e., $\bar{V}_{0}^{{\pi}_{{\bm{1}}/N}}=\frac{1}{N}\sum_{n=1}^{N}V^{{\pi}_{{\bm{1}}/N}}_{0,n}=\frac{1}{S}\sum_{s=1}^{S}V^{{\pi}_{\phi}}_{0,s}=\bar{V}_{0}^{{\pi}_{\phi}}$.
The objective in Eq. 2 is therefore reformulated as $\underset{{\pi}_{\phi}}{\max}\,\bar{V}_{0}^{{\pi}_{\phi}}$, i.e.,
@@
An LP method is provided to solve the count aggregation MDP in Appendix D.2.
@@ 4 COUNT-PROPORTION-BASED DRL
We now consider the situation where the transition dynamics $p^{(N)}_{\phi}$ are unknown and the learner computes the (sub-)optimal policy through trial-and-error interactions with the environment. In Section 4.1, we introduce a count-proportion-based deep RL (CP-DRL) approach. This method incorporates a stochastic policy neural network with fixed-sized inputs and outputs, designed for optimizing resource allocation among stakeholders under constraints with count representation. In Section 4.2, we detail the priority-based sampling procedure used to generate count actions.
@@ 4.1 Stochastic Policy Neural Network
One key property of the count aggregation MDP is that the dimensions of the state space $\mathcal{S}^{(N)}_{f}$ and the action space $\mathcal{A}^{(N)}_{g_{s}}$ are constant and irrespective of the number of sub-MDPs. To further simplify the analysis and eliminate the influence of $N$, we define the count state proportion as $\bar{\bm{x}}=\bm{x}/N$ and the resource proportion constraint for each resource $k$ as $\bar{b}_{k}=b_{k}/(N\max_{a\in\mathcal{A}}d_{k}(a))\in[0,1]$.
@@
Figure 1: CP-based Stochastic Policy Neural Network
@@ 4.2 Priority-based Sampling Procedure
  Apply masking to ${\bm{U}}$ and update the forbidden set by ${\mathcal{F}}\leftarrow\mathcal{F}\cup\{(s,a)\mid u^{p}_{s,a}=0\text{ for }s\in[S]\text{ and }a\in[A]\}$
@@
        Add $(s,a)$ to forbidden set ${\mathcal{F}}\leftarrow\mathcal{F}\cup\{(s,a)\}$
@@ 5 EXPERIMENTAL RESULTS
We apply our methods to the machine replacement problem (Delage and Mannor,, 2010; Akbarzadeh and Mahajan,, 2019), providing a scalable framework for evaluating the CP-DRL approach as problem size and complexity increase. We focus on a single resource ($K=1$) and binary action ($A=2$) for each machine, allowing validation against the Whittle index policy for RMABs (Whittle,, 1988). We applied various DRL algorithms, including Soft Actor Critic (SAC), Twin Delayed DDPG (TD3), and Proximal Policy Optimization (PPO). Among these, PPO algorithm (Schulman et al.,, 2017) consistently delivers the most stable and high-quality performance. We thus choose PPO as the main algorithm for our CP-DRL approach (see Section 4). Our code is provided on GitHub.222 https://github.com/x-tu/GGF-wcMDP.
Machine Replacement Problem   The problem consists of $N$ identical machines following Markovian deterioration rules with $S$ states representing aging stages. The state space $\mathcal{S}^{(N)}$ is a Cartesian product. At each decision stage, actions $\bm{a}$ are applied to all machines under resource constraints, with action $a_{n}$ representing operation (passive action) or replacement (active action). Resource consumption $d_{n}(a_{n})$ is 1 for replacements and 0 for operations, with up to $b$ replacements per time step. The costs range from 0 to 1, transformed to fit the reward representation by multiplying by -1 and adding 1. Machines degrade if not replaced and remain in state $S$ until replaced. Refer to Appendix F.1 for cost structures and transition probabilities. We choose operational and replacement costs across two presets to capture different scenarios (see Appendix F.2 for details): i) Exponential-RCCC and ii) Quadratic-RCCC.
The goal is to find a fair policy that maximizes the GGF score over the expected total discounted mean rewards with count aggregation MDP. In cases like electricity or telecommunication networks (Nadarajah and Cire,, 2024), where equipment is regionally distributed, a fair policy guarantees equitable operations and replacements, thereby preventing frequent failures in specific areas that lead to unsatisfactory and unfair results for certain customers.
Experimental Setup   We designed a series of experiments to test the GGF-optimality, flexibility, scalability, and efficiency of our CP-DRL algorithm. We compare against seven benchmarks, including optimal solutions (OPT) from the GGF-LP model (3) for small instances solved with Gurobi 10.0.3, the Whittle index policy (WIP) for RMABs, and a random (RDM) agent that selects actions randomly at each time step and averages the results over 10 independent runs. Additionally, we implemented a simple DRL baseline, Vanilla-DRL (V-DRL), with a utilitarian objective. The stochastic policy network employs a fully connected neural network that maps the vector $\bm{s}$ to a $N$-dimensional probability vector. We also implemented two heuristics to complement the random agent approach. The oldest-first (OFT) approach selects the machine in the worst state, while the myopic (MYP) selects the machine that maximizes immediate reward. We finally implemented an equal-resources (EQR) approach based on Li and Varakantham, 2022a , which imposes that each machine be replaced once every $N$ steps to ensure an equal distribution of resources.
@@
Experiment 2 (Flexibility)   The fixed-size input-output design allows CP-DRL to leverage multi-task training (MT) with varying machine numbers and resources. We refer to this multi-task extension as CP-DRL(MT). We trained the CP-DRL(MT) with $N\in\{2,3,4,5\}$, randomly switching configurations at the end of each episode over 2000 training episodes. CP-DRL(MT) was evaluated separately, and GGF values for WIP and RDM policies were obtained from 1000 Monte Carlo runs. The numbers following the plus-minus sign ($\pm$) represent the variance across 5 experiments with different random seeds in Table 1 and 2. Variances for WIP and RDM are minimal and omitted, with bold font indicating the best GGF scores at each row excluding optimal values. As shown in Table 1, CP-DRL(MT) consistently achieves scores very close to the OPT values as the number of machines increases from 2 to 4. For the 5-machine case, CP-DRL(MT) shows slightly better performance than the single-task CP-DRL. In Table 2, the single- and multi-task CP-DRL agents show slight variations in performance across different machine numbers. For $N=5$, CP-DRL achieves the best GGF score, slightly outperforming WIP. In both cases, the CP-DRL approach outperforms Vanilla-DRL, the three heuristic methods, and the random agent.
Table 1: GGF Scores (Exponential-RCCC)
$N$ OPT WIP CP-DRL CP-DRL(MT) V-DRL OFT MYP EQR RDM
2 14.19 14.07 14.12 $\pm$ 0.01 14.11 $\pm$ 0.01 13.56 $\pm$ 0.00 5.84 12.59 10.05 9.67
3 14.08 13.75 13.95 $\pm$ 0.02 13.89 $\pm$ 0.14 13.39 $\pm$ 0.00 7.92 12.32 11.67 10.13
4 13.94 13.27 13.64 $\pm$ 0.05 13.59 $\pm$ 0.10 13.04 $\pm$ 0.01 9.02 12.86 12.03 9.74
5 13.77 12.47 12.96 $\pm$ 0.01 13.28 $\pm$ 0.03 12.83 $\pm$ 0.00 10.01 12.08 11.87 8.95
Table 2: GGF Scores (Quadratic-RCCC)
$N$ OPT WIP CP-DRL CP-DRL(MT) V-DRL OFT MYP EQR RDM
2 16.17 16.17 16.14 $\pm$ 0.00 16.14 $\pm$ 0.00 15.36 $\pm$ 0.00 3.11 6.61 9.73 10.15
3 16.10 16.09 16.05 $\pm$ 0.00 16.05 $\pm$ 0.00 15.17 $\pm$ 0.01 6.16 6.63 12.73 11.83
4 16.01 16.01 15.94 $\pm$ 0.00 15.94 $\pm$ 0.00 15.01 $\pm$ 0.00 7.92 6.85 14.02 12.17
5 15.91 15.86 15.87 $\pm$ 0.02 15.86 $\pm$ 0.02 14.73 $\pm$ 0.00 9.25 6.70 14.64 11.98
(a) GGF values for the number of machines $N\in[10,100]$
(b) Time per episode in seconds with a resource ratio $b/N=0.1$
(c) Time per episode in seconds with a resource ratio $b/N=0.5$
@@
Figure 3: (Colored) Scalability and Time Efficiency of CP-DRL.  Subfigures (a) and (b) show the scalability of CP-DRL with a fixed resource ratio of 0.1. Subfigure (a) presents GGF values across different machine counts, with intervals representing the standard deviation over 5 runs. Subfigure (b) and (c) depicts time per episode in seconds for a fixed resource ratio of 0.1 and 0.5, respectively. In all time plots, the green line represents WIP during MC evaluation, the blue line shows CP-DRL during training, and the orange line represents CP-DRL during MC evaluation.
Experiment 3 (Scalability)   We assess CP-DRL scalability by increasing the number of machines while keeping the resource proportion at 0.1 for the Exponential-RCCC instances. We refer to this scaled extension as CP-DRL(SC). We vary the number of machines from 10 to 100 to evaluate CP-DRL performance as the problem size grows. We also use CP-DRL(SC), trained on 10 machines with 1 unit of resource, and scale it to tasks with 20 to 100 machines. Figure 3a shows CP-DRL and CP-DRL(SC) consistently achieve higher GGF values than WIP as machine numbers increase. CP-DRL(SC) delivers results comparable to separately trained CP-DRL, reducing training time while maintaining similar performance. Both WIP and CP-DRL show linear growth in time consumption per episode as machine numbers scale up.
@@
Experiment 4 (Efficiency)   In the GGF-LP model (3), the number of constraints grows exponentially with the number of machines $N$ as $N^{2}+S^{N}$, and the variables increase by $2N+(N+1)\cdot S^{N}$. Using the count dual LP model (20) reduces the model size, but constraints still grow as $\binom{N+S-1}{S-1}$ and variables increase by $\binom{N+S-1}{S-1}\cdot A$. These growth patterns create computational challenges as the problem size increases. A detailed time analysis for the GGF-LP and count dual LP models with $N$ from 2 to 7 is provided in Appendix F.5. In addition to the time per episode for a fixed ratio of 0.1 in Figure 3a, we analyze performance with a 0.5 ratio (Figure 3c) and varying machine proportions, keeping the number of machines fixed at 10. We evaluate CP-DRL over machine proportions from 0.1 to 0.9. The results show that the time per episode increases linearly with the number of machines, while training and evaluation times remain relatively stable. This indicates that the sampling procedure for legal actions is the primary bottleneck. Meanwhile, the resource ratio has minimal impact on computing times.
@@ 6 CONCLUSION
We incorporate the fairness consideration in terms of the generalized Gini function within the weakly coupled Markov decision processes, and define the GGF-WCMDP optimization problem. First, we present an exact method based on linear programming for solving it. We then derive an equivalent problem based on a utilitarian reduction when the WCMDP is symmetric, and show that the set of optimal permutation invariant policy for the utilitarian objective is also optimal for the original GGF-WCMDP problem. We further leverage this result by utilizing a count state representation and introduce a count-proportion-based deep RL approach to devise more efficient and scalable solutions. Our empirical results show that the proposed method using PPO as the RL algorithm consistently achieves high-quality GGF solutions. Moreover, the flexibility provided by the count-proportion approach offers possibilities for scaling up to more complex tasks and context where Whittle index policies are unavailable due to the violation of the indexability property by the sub-MDPs.
@@ Appendix A RELATED WORK
Fairness-aware learning is increasingly integrated into the decision-making ecosystem to accommodate minority interests. However, naively imposing fairness constraints can actually exacerbate inequity (Wen et al.,, 2021) if the feedback effects of decisions are ignored. Many real-world fairness applications are not one-time static decisions (Zhao and Gordon,, 2019) and can thus be better modeled with sequential decision problems, which still remain relatively understudied.
Fairness in resource allocation Fairness has been an important concern in resource allocation problems, where traditional approaches often build upon optimization frameworks, leveraging fairness-constrained optimization (Argyris et al.,, 2022; Chen and Hooker,, 2023), game-theoretic concepts (Namazi and Khodabakhshi,, 2023), or axiomatic principles (Lan and Chiang,, 2011) to derive fair solutions. These models provide practical fair solutions, but do not account for the long-term impact of allocation decisions, which motivates fair-aware sequential decision-making frameworks based on Markov decision processes (MDPs) as discussed next.
@@
Fairness with dynamics There are a few studies investigating fairness-aware sequential decision making without relying on MDPs. For instance, Liu et al., (2018) consider one-step delayed feedback effects, Creager et al., (2020) propose causal modeling of dynamical systems to address fairness, and Zhang et al., (2019) construct a user participation dynamics model where individuals respond to perceived decisions by leaving the system uniformly at random. These studies extend the fairness definition in temporally extended decision-making settings, but do not take feedback and learning into account that the system may fail to adapt to changing conditions. Alamdari et al., (2023) address this gap by introducing multi-stakeholder fairness as non-Markovian sequential decision making and developing a Q-learning based algorithm with counterfactual experiences to enhance sample-efficient fair policy learning.
@@
Fairness in Markov decision processesZhang et al., (2020) consider how algorithmic decisions impact the evolution of feature space of the underlying population modeled as MDPs but is limited to binary decisions. Ghalme et al., (2022) study a fair resource allocation problem in the average MDP setting and proposes an approximate algorithm to compute the policy with sample complexity bounds. However, their definition of fairness is restricted to the minimum visitation frequency across all states, potentially resulting in unbalanced rewards among sub-MDPs. Wen et al., (2021) develop fair decision-making policies in discounted MDPs, but the performance guarantees are achieved only under a loose condition. In contrast, our work takes into account a more comprehensive definition of fairness. Segal et al., (2023) investigate the impact of societal bias dynamics on long-term fairness and the interplay between utility and fairness under various optimization parameters. Additionally, Hassanzadeh et al., (2023) address a fair resource allocation problem similar to our work but in continuous state and action space. They define fairness to the agents considering all their allocations over the horizon under the Nash Social Welfare objective in hindsight.
@@
Fairness in reinforcement learningJabbari et al., (2017) initiate the meritocratic fairness notion from the multi-arm bandits setting to the reinforcement learning (RL) setting. Later, fairness consideration has been integrated in RL to achieve fair solutions in different domains, including a fair vaccine allocation policy that equalizes outcomes in the population (Atwood et al.,, 2019), balancing between fairness and accuracy for interactive user recommendation (Liu et al.,, 2020; Ge et al.,, 2022), and fair IoT that continuously monitors the human state and changes in the environment to adapt its behavior accordingly (Elmalaki,, 2021). However, most work focuses on the impartiality aspect of fairness. Jiang and Lu, (2019) investigate multi-agent RL where fairness is defined over agents and encoded with a different welfare function, but the focus is on learning decentralized policies in a distributed way. We refer readers to two literature review papers by Gajane et al., (2022) and Reuel and Ma, (2024) on fairness considerations in RL, which provide comprehensive insights into current trends, challenges, and methodologies in the field.